OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. -4 + yi = x + 3i

OpenStudy (radar):

-4 + yi = x + 3i -x + yi = 4 + 3i -x = 4, x=-4 y = 3 I may be thinking wrong but won't that work.